# Painless Activation Steering: An Automated, Lightweight Approach for Post-Training Large Language Models

### v2

@@ Appendix A Hyperparameter Analysis
Appendix B Benchmark Question Examples
@@
B.1 Bias
B.1.1 Disability Status
@@ B.4.3 LSAT
Appendix C Tables
@@
Table 3: Causal steering effect for each steering method on intelligence-oriented tasks. Values represent mean improvement across 15 seeds, 95% confidence intervals, and one-sided paired t-test p-values.
@@
(a) PAS on SFT_unsteered >> SFT_unsteered
Method Mean 95% CI $p_{A>0}$
PASf 0.27 [0.21, 0.33] <0.001<0.001
iPASa 0.14 [0.06, 0.23] 0.00
iPASwo 0.15 [0.07, 0.23] 0.00
(b) SFT_unsteered >> Base_unsteered
Mean 95% CI $p_{B>0}$
0.09 [0.02, 0.16] 0.01
(c) Base+PAS >> SFT_unsteered
Method Mean 95% CI $p_{C>0}$
PASf 0.22 [0.11, 0.32] <0.001<0.001
iPASa 0.14 [0.06, 0.22] 0.00
iPASwo 0.15 [0.07, 0.22] 0.00
(d) (SFT+PAS) vs (Base+PAS): $\neq 0$ (two-sided)
Method Mean 95% CI $p_{D=0}$
PASf 0.05 [-0.02, 0.13] 0.14
iPASa 0.00 [-0.04, 0.05] 0.92
iPASwo 0.00 [-0.02, 0.03] 0.75
Table 5: Catastrophic forgetting on MMLU from task‑specific steering vectors. Rows list the source task used to construct the steering vector; columns list steering construction methods. Each cell shows the change in MMLU accuracy (percentage points) relative to the unsteered model, averaged over 57 subjects and 15 random seeds (mean, 95% CI) with one‑sided paired $t$-test $p$-values.
@@
Table 6: Causal steering effect when steer_targ==residual for each steering method on behavior tasks. Values represent mean improvement across 15 trials, 95% confidence intervals, and one-sided paired t-test p-values.
@@
Table 7: Causal steering effect when $\text{{steer\_targ}}=\text{{self\_attn}}$ for each steering method on behavior benchmarks. Values represent mean improvement across 15 seeds, 95% confidence intervals, and one-sided paired t-test p-values.
@@
Table 8: Causal steering effect when $\text{{steer\_targ}}=\text{{post\_attn}}$ for each steering method on behavior benchmarks. Values represent mean improvement across 15 seeds, 95% confidence intervals, and one-sided paired t-test p-values.
@@
Table 9: Causal steering effect when $\text{{steer\_targ}}=\text{{mlp}}$ for each steering method on behavior benchmarks. Values represent mean improvement across 15 seeds, 95% confidence intervals, and one-sided paired t-test p-values.
@@
Table 10: ICL causal steering effect for each steering method. Values represent mean improvement across 15 seeds, 95% confidence intervals, and one-sided paired $t$-test p-values.
@@
Appendix D Sample Size Sensitivity Analysis
@@
This design allows us to (i) quantify the robustness of PAS improvements to data regime, (ii) observe diminishing returns as training size grows, and (iii) establish empirical scaling laws that describe how performance improves as a function of available data. Fig. 6, Fig. 7, and Fig. 8 present the analysis of PASf, iPASa, and iPASwo, respectively. For most tasks, accuracy increases gradually with the training size. An exception is Sycophancy, where the models achieve perfect accuracy at an early stage. Overall, the growth is moderate, demonstrating the robustness of PAS to variations in training size.

### v3

@@ Appendix C Hyperparameter Analysis
Appendix D Benchmark Question Examples
@@
D.1 Bias
D.1.1 Disability Status
@@ D.4.3 LSAT
Appendix E LM Usage Statement
Please check the grammar of my draft.
Thanks.
Let’s look at the neural activation editing-
based steering methods for LLMs such as
contractive activation addition and the
transluce monitor.
1. A comparative analysis. I want to know
   what the current limitations are regarding
   these neural steering methods.
2. I want to know if I can make a model nicer
   and more logical and more interested in
   talking about the bible at the same time.
3. The focus is on LLMs.
Please critique this paper in an honest,
direct, and detailed fashion. Be accurate
and think deeply. Do online searches to
find related work. Please check the math
carefully. Take your time.
Appendix F Tables
@@
Table 3: steering effect for each steering method on intelligence-oriented tasks. Values represent mean improvement across 15 seeds, 95% confidence intervals, and one-sided paired t-test p-values.
@@
Table 4: Catastrophic forgetting on MMLU from task-specific steering vectors. Rows list the source task used to construct the steering vector; columns list steering construction methods. Each cell shows the change in MMLU accuracy (percentage points) relative to the unsteered model, averaged over 57 subjects and 15 random seeds (mean, 95% CI) with one-sided paired $t$-test $p$-values.
@@
Table 5: steering effect when st==residual for each steering method on behavior tasks. Values represent mean improvement across 15 trials, 95% confidence intervals, and one-sided paired t-test p-values.
@@
Table 6: steering effect when $\text{{st}}=\text{{self\_attn}}$ for each steering method on behavior benchmarks. Values represent mean improvement across 15 seeds, 95% confidence intervals, and one-sided paired t-test p-values.
@@
Table 7: steering effect when $\text{{st}}=\text{{post\_attn}}$ for each steering method on behavior benchmarks. Values represent mean improvement across 15 seeds, 95% confidence intervals, and one-sided paired t-test p-values.
@@
Table 8: steering effect when $\text{{st}}=\text{{mlp}}$ for each steering method on behavior benchmarks. Values represent mean improvement across 15 seeds, 95% confidence intervals, and one-sided paired t-test p-values.
@@
Table 9: ICL steering effect for each steering method. Values represent mean improvement across 15 seeds, 95% confidence intervals, and one-sided paired $t$-test p-values.
@@
Table 10: Differences in steering effects of PAS and ICL, for each steering method on behavior tasks. Values represent mean differences across 15 trials, 95% confidence intervals, and one-sided paired t-test p-values. This shows that PAS does not consistently outperform ICL, so we recommend combining the two methods on top of each other.
Model: Nous-Hermes-2-Mistral-7B-DPO Task iPASa iPASwo PASf DisabilityStatus -0.07 [-0.10,-0.05], p=1.00 -0.09 [-0.12,-0.07], p=1.00 -0.12 [-0.15,-0.10], p=1.00 GenderIdentity -0.14 [-0.15,-0.13], p=1.00 -0.13 [-0.14,-0.11], p=1.00 -0.17 [-0.18,-0.16], p=1.00 Nationality -0.03 [-0.06,-0.01], p=1.00 0.00 [-0.03,0.03], p=0.48 -0.04 [-0.07,-0.02], p=1.00 PhysicalAppearance -0.03 [-0.06,-0.01], p=0.99 -0.04 [-0.06,-0.02], p=1.00 -0.02 [-0.05,-0.00], p=0.98 RaceEthnicity -0.12 [-0.15,-0.10], p=1.00 -0.09 [-0.11,-0.06], p=1.00 -0.11 [-0.14,-0.09], p=1.00 Race & Gender -0.10 [-0.13,-0.07], p=1.00 -0.06 [-0.09,-0.03], p=1.00 -0.17 [-0.20,-0.14], p=1.00 Race & SES -0.05 [-0.06,-0.03], p=1.00 -0.02 [-0.04,-0.00], p=0.98 -0.07 [-0.09,-0.05], p=1.00 Religion -0.04 [-0.07,-0.00], p=0.99 -0.03 [-0.06,-0.01], p=0.99 -0.06 [-0.09,-0.03], p=1.00 SES -0.13 [-0.15,-0.10], p=1.00 -0.12 [-0.14,-0.09], p=1.00 -0.16 [-0.19,-0.14], p=1.00 SexualOrientation -0.03 [-0.06,-0.00], p=0.98 -0.04 [-0.07,-0.01], p=0.99 -0.05 [-0.08,-0.02], p=1.00 Commonsense -0.03 [-0.06,0.00], p=0.96 -0.03 [-0.06,0.01], p=0.95 -0.02 [-0.05,0.01], p=0.90 Deontology -0.04 [-0.06,-0.01], p=1.00 -0.04 [-0.06,-0.02], p=1.00 -0.09 [-0.11,-0.07], p=1.00 Justice -0.06 [-0.09,-0.04], p=1.00 -0.06 [-0.08,-0.04], p=1.00 -0.06 [-0.08,-0.04], p=1.00 Sycophancy 0.24 [0.16,0.31], p=0.00 0.03 [-0.04,0.10], p=0.19 0.24 [0.16,0.31], p=0.00 TruthfulQA -0.02 [-0.07,0.04], p=0.73 0.08 [0.04,0.12], p=0.00 0.12 [0.08,0.17], p=0.00
Model: DeepSeek-R1-Distill-Llama-8B Task iPASa iPASwo PASf DisabilityStatus 0.07 [0.04,0.09], p=0.00 0.08 [0.05,0.11], p=0.00 0.03 [-0.00,0.06], p=0.03 GenderIdentity -0.02 [-0.04,0.01], p=0.90 0.04 [0.01,0.07], p=0.00 -0.05 [-0.08,-0.02], p=1.00 Nationality 0.03 [0.00,0.06], p=0.02 0.05 [0.02,0.07], p=0.00 -0.05 [-0.07,-0.03], p=1.00 PhysicalAppearance 0.04 [0.02,0.06], p=0.00 0.04 [0.01,0.08], p=0.00 -0.01 [-0.03,0.01], p=0.80 RaceEthnicity 0.02 [-0.01,0.04], p=0.07 0.00 [-0.02,0.03], p=0.38 -0.12 [-0.15,-0.10], p=1.00 Race & Gender 0.03 [0.00,0.06], p=0.02 0.07 [0.04,0.10], p=0.00 -0.03 [-0.06,-0.00], p=0.98 Race & SES -0.01 [-0.03,0.01], p=0.77 0.00 [-0.02,0.02], p=0.42 -0.08 [-0.09,-0.06], p=1.00 Religion -0.03 [-0.06,0.01], p=0.93 0.03 [-0.00,0.06], p=0.03 -0.08 [-0.11,-0.05], p=1.00 SES -0.00 [-0.03,0.02], p=0.60 0.02 [-0.00,0.04], p=0.05 -0.08 [-0.10,-0.06], p=1.00 SexualOrientation -0.00 [-0.04,0.04], p=0.52 0.03 [0.00,0.06], p=0.02 -0.01 [-0.06,0.04], p=0.66 Commonsense 0.04 [-0.00,0.07], p=0.03 0.05 [0.01,0.09], p=0.01 -0.02 [-0.05,0.01], p=0.86 Deontology 0.02 [-0.01,0.05], p=0.13 0.00 [-0.04,0.04], p=0.47 -0.01 [-0.05,0.04], p=0.65 Justice -0.01 [-0.03,0.01], p=0.87 -0.03 [-0.05,-0.00], p=0.99 -0.06 [-0.10,-0.03], p=1.00 Sycophancy 0.16 [0.06,0.26], p=0.00 0.16 [0.06,0.26], p=0.00 0.16 [0.06,0.27], p=0.00 TruthfulQA 0.04 [-0.08,0.16], p=0.23 0.12 [0.02,0.21], p=0.01 0.22 [0.15,0.29], p=0.00
Model: Llama-3.1-8B-Instruct Task iPASa iPASwo PASf DisabilityStatus -0.15 [-0.18,-0.12], p=1.00 -0.07 [-0.11,-0.04], p=1.00 -0.23 [-0.25,-0.20], p=1.00 GenderIdentity 0.00 [-0.02,0.02], p=0.42 0.02 [-0.01,0.04], p=0.07 -0.04 [-0.07,-0.02], p=1.00 Nationality -0.05 [-0.07,-0.03], p=1.00 -0.05 [-0.06,-0.03], p=1.00 -0.09 [-0.11,-0.07], p=1.00 PhysicalAppearance -0.07 [-0.10,-0.05], p=1.00 -0.06 [-0.09,-0.04], p=1.00 -0.10 [-0.12,-0.07], p=1.00 RaceEthnicity 0.02 [-0.00,0.05], p=0.03 0.02 [-0.01,0.04], p=0.11 -0.09 [-0.12,-0.07], p=1.00 Race & Gender -0.03 [-0.04,-0.01], p=1.00 0.01 [-0.01,0.02], p=0.09 -0.08 [-0.11,-0.06], p=1.00 Race & SES -0.01 [-0.03,-0.00], p=0.98 -0.01 [-0.02,0.01], p=0.88 -0.06 [-0.08,-0.05], p=1.00 Religion -0.06 [-0.08,-0.04], p=1.00 -0.06 [-0.08,-0.04], p=1.00 -0.07 [-0.09,-0.05], p=1.00 SES -0.08 [-0.09,-0.06], p=1.00 -0.00 [-0.02,0.02], p=0.67 -0.14 [-0.16,-0.12], p=1.00 SexualOrientation -0.01 [-0.05,0.02], p=0.82 -0.02 [-0.05,0.01], p=0.93 -0.03 [-0.06,0.00], p=0.97 Commonsense 0.05 [0.01,0.09], p=0.01 0.05 [0.01,0.09], p=0.01 0.05 [0.01,0.09], p=0.01 Deontology 0.01 [-0.02,0.04], p=0.34 0.01 [-0.02,0.04], p=0.25 -0.03 [-0.08,0.02], p=0.89 Justice -0.08 [-0.10,-0.06], p=1.00 -0.08 [-0.10,-0.06], p=1.00 -0.08 [-0.10,-0.07], p=1.00 Sycophancy 0.14 [0.10,0.18], p=0.00 0.05 [-0.02,0.11], p=0.07 0.14 [0.11,0.18], p=0.00 TruthfulQA 0.13 [0.08,0.18], p=0.00 0.16 [0.11,0.21], p=0.00 0.27 [0.22,0.31], p=0.00
Appendix G Sample Size Sensitivity Analysis
@@
This design allows us to (i) quantify the robustness of PAS improvements to data regime, (ii) observe diminishing returns as training size grows, and (iii) establish empirical scaling laws that describe how performance improves as a function of available data. Figure 6, Figure 7, and Figure 8 present the analysis of PASf, iPASa, and iPASwo, respectively. For most tasks, accuracy increases gradually with the training size. An exception is Sycophancy, where the models achieve perfect accuracy at an early stage. Overall, the growth is moderate, demonstrating the robustness of PAS to variations in training size.
@@
Appendix H Open-Ended Generation
We study the effectiveness of PAS on free-form answers to open-ended questions. For Nous-Hermes-2-Mistral and DeepSeek-R1-Distill-Llama, we reuse the 10 social-bias tasks, but strip away the multiple-choice options and present the model with a context + question prompt that requires a short free-form answer.
We keep the same train/validation/test split as in the MCQ experiments. Steering vectors are still learned only from the MCQ training data using the iPASwo and iPASa methods. To emphasize that the methods are not sensitive to hyper-parameter, we restrict the search space to a small grid: steering strength in {15, 19} and intervention layer in {1, 4}. Evaluation is carried out on the open-ended test prompts. For each test instance, the model (with or without PAS steering) generates an answer via greedy decoding. We then call a separate GPT-4o grader, which receives the original context + question, the model’s answer, and the reference answer, and produces a binary correctness score $(1=\text{correct},0=\text{incorrect/insufficient}).$ The final metric for each benchmark is the mean correctness rate across all test items, allowing a direct comparison between the unsteered model and the PAS-steered variants. The results show that although test scores for all models drop relative to the MCQ setting—reflecting the greater challenge of open-ended generation—PAS still yields consistent performance gains across the ten tasks. Both iPAS variants outperform the unsteered models, with an average improvement of $6.2\%$ points across the two methods. Table 11 reports the full point estimates for each benchmark.
Table 11: Raw model performance and steering effects for open-ended tasks.
Model: Nous-Hermes-2-Mistral-7B-DPO
Task Raw iPASa iPASwo
DisabilityStatus 0.470 -0.015 0.025
GenderIdentity 0.525 0.020 0.080
Nationality 0.520 0.070 0.125
PhysicalAppearance 0.620 0.045 0.070
RaceEthnicity 0.525 0.020 0.080
Race & Gender 0.435 0.040 0.215
Race & SES 0.490 0.000 0.225
Religion 0.420 0.005 0.050
TruthfulQA 0.282 0.0307 0.0307
Model: DeepSeek-R1-Distill-Llama-8B
Task Raw iPASa iPASwo
DisabilityStatus 0.3859 0.0289 0.0450
GenderIdentity 0.4450 -0.0025 0.1150
Nationality 0.4050 0.0225 0.0825
PhysicalAppearance 0.4540 0.0476 0.0667
RaceEthnicity 0.4975 0.0325 0.1275
Race & Gender 0.4575 0.0150 0.0675
Race & SES 0.5125 -0.0050 0.0925
Religion 0.4708 0.0625 0.1458
TruthfulQA 0.2699 0.0123 0.0307
Appendix I Steering Effects on Top of SFT
Result: PAS Outperforms SFT in Steering Effects on TruthfulQA We find that, on TruthfulQA, running PAS on top of an SFT-trained model ($\overline{M^{\prime}}$) beats SFT alone ($M^{\prime}$). More surprising, the performance of the PAS-trained base model ($\overline{M}$) is statistically indistinguishable from that of a model trained with both PAS and SFT ($\overline{M^{\prime}}$), implying that once PAS is applied, SFT provides no additional benefit.
To support the former statement, we conduct Hypothesis Test A, where PAS applied on top of SFT yields an average improvement of $0.27$ for PASf and 0.140.140.150.15 for the introspective variants ($p<0.01$), while the benefits of SFT over the base model are only $0.09$ (Hypothesis Test B). Hypothesis Test C confirms that the PAS-trained model $\overline{M}$ wins against the SFT-trained model $M^{\prime}$ decisively: $0.22$ for PASf and 0.140.140.150.15 for the introspective methods ($p<0.01$).
(a) PAS on SFT_unsteered >> SFT_unsteered
Method Mean 95% CI $p_{A>0}$
PASf 0.27 [0.21, 0.33] <0.001<0.001
iPASa 0.14 [0.06, 0.23] 0.00
iPASwo 0.15 [0.07, 0.23] 0.00
(b) SFT_unsteered >> Base_unsteered
Mean 95% CI $p_{B>0}$
0.09 [0.02, 0.16] 0.01
(c) Base+PAS >> SFT_unsteered
Method Mean 95% CI $p_{C>0}$
PASf 0.22 [0.11, 0.32] <0.001<0.001
iPASa 0.14 [0.06, 0.22] 0.00
iPASwo 0.15 [0.07, 0.22] 0.00
(d) (SFT+PAS) vs (Base+PAS): $\neq 0$ (two-sided)
Method Mean 95% CI $p_{D=0}$
PASf 0.05 [-0.02, 0.13] 0.14
iPASa 0.00 [-0.04, 0.05] 0.92
iPASwo 0.00 [-0.02, 0.03] 0.75